OpenStudy (anonymous):

if v = At^3 - Bt^4 what are the dimensions of A? what are the dimensions of B?

OpenStudy (anonymous):

dimention for v must be length/time if it is velocity or speed so \[v = t ^{3} (A - Bt)\] so dimentions for (A - Bt) must be length/time^3 and thus A has dimentions length/time^3 and B length/time^4

OpenStudy (anonymous):

As there are two unknowns in one equation, A and B have to be expressed with respect to each other. v = d/t Therefore the equation can be written as d/t = t^3(A-Bt) d/t^4 = A-Bt So A = d/t^4 + Bt and B = t^-1(A-dt^-4)